# Causal Inference: A Tale of Three Frameworks

### v1

@@ 2.3 Directed Acyclic Graphs
The use of DAGs for causal inference stems from foundational work by Spirtes et al., (2000) and Pearl, (1995), building on earlier ideas from Sewall Wright’s path diagrams (Wright,, 1921). In this framework, a causal system is represented by a DAG, where nodes correspond to variables and directed edges encode possible direct causal effects. In particular, the absence of an edge asserts that no direct causal effect is present. In the following, we introduce the DAG model from three perspectives: purely graphical, statistical/probabilistic, and causal.
@@
Some authors refer to the subpaths in 1. as non-colliders. If all paths from nodes in $A$ to nodes in $B$ are blocked by $C$, then $A$ and $B$ are said to be d-separated by $C$ (the d denotes directional (Pearl et al.,, 2016, p. 46)). Equivalently, d-separation can be defined using the notion of “moral graph” (Lauritzen et al.,, 1990). Specifically, consider the subgraph induced by the ancestors of $A\cup B\cup C$. For each pair of non-adjacent nodes that share a common child, add an undirected edge between them (thus “marrying” the parent variables). Then, remove the directions of all edges to obtain an undirected moral graph. We say $A$ and $B$ are d-separated by $C$ in the original DAG if and only if every path between $A$ and $B$ is intercepted by a node in $C$ in the resulting moral graph.
Unlike the concepts of parents and descendants, d-separation is a relatively modern development (Pearl and Paz,, 1986; Pearl,, 1988; Geiger et al.,, 1990; Pearl,, 1995). It enables one to read off conditional independencies from directed graphs in probabilistic DAG models, to which we now turn.
@@
In probabilistic or statistical modeling, a DAG is sometimes referred to as a Bayesian network (Pearl,, 1985), and each node in a DAG represents a random variable. Here the DAG is used to encode the conditional independence structure of a joint distribution over a set of random variables $(V_{1},\ldots,V_{p})$. Specifically, let $P$ be a joint distribution over $(V_{1},\ldots,V_{p})$, and let $\mathcal{G}$ be a DAG. We say that $P$ is Markovian with respect to $\mathcal{G}$ if it satisfies the following properties.
@@
These Markov properties are equivalent if the joint distribution of $V$ has a density with respect to a product measure (Lauritzen,, 1996, Theorem 6.22).
@@
Different graphs may encode the same set of conditional independence relationships. For example, the graphs $A\rightarrow B\rightarrow C$ and $A\leftarrow B\rightarrow C$ both encode the assumption that $A\perp\!\!\!\perp C\mid B$. In this case, these graphs are called Markov equivalent. Graphs that are Markov equivalent cannot be distinguished based solely on conditional independencies in observational data. However, they may be distinguished in the presence of additional assumptions about the joint distribution, such as constraints on the functional form of the conditional distributions. For detailed discussions, see Example 3 in Section 4.2 and Peters et al., (2017, §4 & §7).
@@
It can be shown that the set of distributions unfaithful to a graph $\mathcal{G}$ has Lebesgue measure zero (Spirtes et al.,, 2000, Theorem 3.2); see also Uhler et al., (2013). This result, however, comes with an important caveat: there exist sequences of faithful distributions that converge arbitrarily close to unfaithful distributions. A more robust alternative is the concept of $\lambda$-strong faithfulness, which requires that the dependence strength (e.g., partial correlation) between d-connected variables exceeds some threshold $\lambda>0$. Unlike exact faithfulness, the set of distributions that are not $\lambda$-strong faithful may occupy a substantial volume in the space of distributions (Uhler et al.,, 2013).
@@
Strictly speaking, the phrase “sample uniformly” is not well defined without first specifying a parameterization, a point related to Borel’s paradox. Conditional distributions on measure zero sets can depend on the chosen parameterization (Billingsley,, 1995, p. 441, Problem 33.1); see also Wang, (2022). In particular, subsets such as $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ are lower dimensional manifolds within the probability simplex, so there is no canonical uniform measure on them. For $\mathcal{S}_{1}$ and the associated DAG $B\rightarrow D\leftarrow A$, we make this precise by drawing each conditional probability table independently from a $\text{Dirichlet}(1,\ldots,1)$ prior and combining them using the DAG factorization, under which the set of unfaithful distributions has Lebesgue measure zero. For $\mathcal{S}_{2}$, we describe in Appendix A1 a sampling scheme under which a distribution satisfying $A\perp B$ and $A\perp D$ almost surely does not satisfy $A\perp(B,D)$.
@@
So far, we have described how DAGs encode statistical conditional independencies via the factorization of the observed data distribution. To endow a DAG with causal meaning, we must go beyond this statistical interpretation and commit to assumptions about how the system behaves under external interventions (Pearl,, 2009, Definition 1.3.1); see also Peters et al., (2017, Definition 6.32).
@@
Eqn. (6) was originally introduced by Robins, (1986) and termed the g-computation algorithm formula or simply the g-formula. It was subsequently independently re-discovered and named the manipulation theorem (Spirtes et al.,, 2000) and the truncation formula (Pearl,, 2009).
When a causal DAG is postulated, it is often assumed that the set of variables $V$ is causally sufficient, in the sense that there are no unmeasured common causes $C\notin V$ that cause two or more variables in $V$ (Spirtes,, 2010). This assumption may be slightly relaxed to the notion of interventional sufficiency, which is sufficient to ensure that the truncation formula (6) holds; see Peters et al., (2017, § 9.1) for a detailed discussion. When unmeasured common causes do exist between variables, it is conventional to represent them by adding bidirected edges ($\leftrightarrow$) between the affected nodes, resulting in an Acyclic Directed Mixed Graph (ADMG) (Richardson,, 2003). Formally, such graphs can be obtained as latent projections of DAGs with unobserved variables (Verma and Pearl,, 1991; Richardson and Spirtes,, 2002).
@@
When all variables in the DAG are observed, it is sufficient to include all the parents of $A$ in the adjustment set $L$, in which case the average causal effect (ACE) is identified by eqn. (2) (Pearl,, 2009, Theorem 3.2.2). More generally, even if some variables in the DAG are unobserved, the following backdoor criterion provides a graphical condition under which causal effects can still be identified from observational data.
@@
Under this criterion, the ACE is still identified by eqn. (2) (Pearl,, 2009, Theorem 3.3.2).
@@
Although Pearl, (2009, § 3.2.1, eqn. (3.2)) interprets the causal semantics of a DAG using the NPSEM-IE framework (see also Section 3.2), the proof of the backdoor criterion (Pearl,, 2009, Theorem 3.2.2 and § 11.3.3) relies only on the g-formula (6), and therefore holds under the causal DAG model.
The backdoor criterion can be operationalized using the following simple rules (Pearl,, 2009, § 11.3.1), such that after conditioning on $L$,
@@ 2. Rule 2:
The first rule of the backdoor criterion suggests that adjusting for post-treatment variables may introduce bias. To illustrate, consider the causal DAG in Figure 2(a), where the treatment $A$ has two descendants, $S$ and $Z$, in addition to the primary variables $A$ and $Y$. Adjusting for $S$ clearly violates Rule 2 above. Now consider adjusting for $Z$. On the surface, $Z$ does not lie on any path from $A$ to $Y$, and $S$ is not a collider on the path $A\rightarrow S\rightarrow Y$. However, every observed variable has an associated error term, and in panel (a) we explicitly display $\varepsilon_{S}$, the error for $S$, since it plays a key role here. Conditioning on $Z$ induces a dependence between $A$ and $\varepsilon_{S}$, which in turn opens the non-causal path $A\leftrightarrow\varepsilon_{S}\to S\to Y$. Equivalently, this corresponds to the undirected path $A-\varepsilon_{S}-S-Y$ in the induced moral graph. Another way to see this is by marginalizing out $S$ and examining the resulting causal DAG in Figure 2(b). In this graph, it becomes clear that $Z$ is a collider on the non-causal path $A\rightarrow Z\leftarrow\varepsilon_{S}\rightarrow S$, and hence adjusting for $Z$ may introduce collider bias into the causal effect estimate (Pearl,, 2009, § 11.3.1).
@@
When the backdoor criterion fails to apply directly, identification may still be possible through Pearl’s do-calculus (Pearl,, 1995), a set of three transformation rules that enable the conversion of interventional expressions (e.g., $P_{\text{do}(a)}(Y)$) into estimable quantities from the observational distribution. While the backdoor criterion suffices for simple DAGs, complete graphical criteria for covariate adjustment in ancestral and partially directed graphs have been developed (Perković et al.,, 2018). In fact, any identifiable interventional distribution under a causal DAG model that may involve unmeasured nodes can be derived through recursive application of the do-calculus rules, as formalized by the ID algorithm (Tian and Pearl,, 2002; Huang and Valtorta,, 2006; Shpitser and Pearl,, 2006). See Tian and Shpitser, (2010) for a review of these results.
It is important to note that this result pertains to identifiability under the causal DAG model alone. Given additional assumptions, identification may still be possible even when the ID algorithm fails to apply. For instance, one can show that the ACE is not identifiable under the instrumental variable model alone (Balke and Pearl,, 1997). However, it can still be identified under additional effect homogeneity assumptions (Wang and Tchetgen Tchetgen,, 2018, Theorem 1); see also Hartwig et al., (2023, Theorem 1) and Dong et al., (2025, Theorem 1).
@@
There are more sophisticated algorithms that also identify counterfactual quantities, such as those discussed in Remark 3 (Shpitser and Pearl,, 2007, 2008). These algorithms, unlike the ID algorithm mentioned above, rely on the NPSEM-IE framework.
@@ 3 Translating Between the Three Causal Frameworks
The question of which causal framework is best suited for practical applications has been the subject of extensive debate in the causal inference literature (e.g. Pearl,, 1995; Rubin,, 2004; Lauritzen,, 2004; Dawid,, 2015; Richardson and Robins,, 2023); see also Pearl, (2009, p. 106). While each approach, including potential outcomes, structural causal models, and graphical models, offers unique strengths, efforts have been made to reconcile and translate among them. Much like natural languages, certain causal concepts are more naturally expressed in one framework than another, making such translations both valuable when possible but, at times, inherently limited. In this section, we review key attempts to bridge these perspectives. Detailed comparisons of these frameworks are deferred to Section 4.
@@ 3.1 The Potential Outcomes and NPSEM-IE Frameworks
Under an NPSEM-IE, each endogenous variable $V_{j}$ is determined by a deterministic function of its parents and an exogenous noise variable $\varepsilon_{j}$, with the collection $\{\varepsilon_{j}\}$ assumed to be jointly independent. This structural system implicitly defines a rich set of potential outcomes. For example, consider the NPSEM in (4). In this case, the potential outcome $Y(l,a)$ is obtained by intervening to set $L=l$, $A=a$, and evaluating the outcome function at the realized value of the noise term $\varepsilon_{Y}$ (Pearl,, 2009, §3.6.3, §7.1):
@@
see also Strotz and Wold, (1960). In fact, researchers have proposed expressing NPSEMs using potential outcomes notation to clearly distinguish them from standard statistical regression models (e.g. Imbens,, 2014; Richardson and Robins,, 2014, 2023). For example, the NPSEM in (4) can be represented as
@@
In the NPSEM defined in (3), the set of input variables $U_{j}$ for a given node $V_{j}$ often does not include all variables that precede $V_{j}$ in the causal ordering. The omission of these variables reflects an implicit exclusion restriction on the corresponding potential outcomes (Pearl,, 2009, p. 101, eqn. (3.55)):
@@
Eqn. (13) is often known as the exclusion restriction in the IV literature (Angrist et al.,, 1996).
@@
Robins and Richardson, (2011); Richardson and Robins, (2013) formalize the converse direction by showing how one can construct an NPSEM-IE from a given system of potential outcomes using what they term one-step-ahead counterfactuals, i.e., the potential outcome of a variable under interventions on all of its parents (Robins,, 1986). This “counterfactual interpretation of the error terms in structural equation models” was also noted in Pearl, (2009, §7.4.5) and referred to as the “canonical representation of structural causal models” by Peters et al., (2017, §3.4).
@@ 3.2 NPSEM-IE and Causal DAG
The first claim in Proposition 1 follows from Pearl, (2009, Theorem 1.4.1); see also Peters et al., (2017, Proposition 6.31). The second claim holds because, under the autonomy assumption, the structural equation for the intervened variable is replaced while all others remain unchanged; see also Strotz and Wold, (1960, §2), Pearl, (2009, p. 72 eqn. (3.10)) and Peters et al., (2017, p. 109, eqn. (6.7)).
@@
In fact, a stronger claim holds: the Single-World Intervention Graph (SWIG) model (see §3.3 below) with interventions on all variables also implies a causal DAG model (Richardson and Robins,, 2023, Theorem 11). This is stronger because the NPSEM-IE is a strict submodel of the SWIG model associated with the same DAG. The converse is not necessarily true. For example, the SWIG in Figure 3 implies the so-called extended g-formula that
@@
The Functional Representation Lemma (El Gamal and Kim,, 2011, p. 626, Appendix B) ensures that any random variable with a known conditional distribution can be expressed as a measurable function of its parents and an independent noise variable. A concrete construction is provided in Peters et al., (2017, Appendix C.9).
@@
where each $f_{j}$ is a measurable function, each $\varepsilon_{j}\sim\text{Uniform}[0,1]$, and the noise variables $\varepsilon_{1},\dots,\varepsilon_{p}$ are jointly independent. The NPSEM (16) with independent errors is compatible with both the causal DAG $\mathcal{G}$ and the joint distribution $P$. A similar construction for the discrete case can be found in Druzdzel and Simon, (1993).

### v2

@@ 2.3 Directed Acyclic Graphs
The use of DAGs for causal inference stems from foundational work by Spirtes et al. (2000) and Pearl (1995), building on earlier ideas from Sewall Wright’s path diagrams (Wright, 1921). In this framework, a causal system is represented by a DAG, where nodes correspond to variables and directed edges encode possible direct causal effects. In particular, the absence of an edge asserts that no direct causal effect is present. In the following, we introduce the DAG model from three perspectives: purely graphical, statistical/probabilistic, and causal.
@@
Some authors refer to the subpaths in 1. as non-colliders. If all paths from nodes in $A$ to nodes in $B$ are blocked by $C$, then $A$ and $B$ are said to be d-separated by $C$ (the d denotes directional (Pearl et al., 2016, p. 46)). Equivalently, d-separation can be defined using the notion of “moral graph” (Lauritzen et al., 1990). Specifically, consider the subgraph induced by the ancestors of $A\cup B\cup C$. For each pair of non-adjacent nodes that share a common child, add an undirected edge between them (thus “marrying” the parent variables). Then, remove the directions of all edges to obtain an undirected moral graph. We say $A$ and $B$ are d-separated by $C$ in the original DAG if and only if every path between $A$ and $B$ is intercepted by a node in $C$ in the resulting moral graph.
Unlike the concepts of parents and descendants, d-separation is a relatively modern development (Pearl and Paz, 1986; Pearl, 1988; Geiger et al., 1990; Pearl, 1995). It enables one to read off conditional independencies from directed graphs in probabilistic DAG models, to which we now turn.
@@
In probabilistic or statistical modeling, a DAG is sometimes referred to as a Bayesian network (Pearl, 1985), and each node in a DAG represents a random variable. Here the DAG is used to encode the conditional independence structure of a joint distribution over a set of random variables $(V_{1},\ldots,V_{p})$. Specifically, let $P$ be a joint distribution over $(V_{1},\ldots,V_{p})$, and let $\mathcal{G}$ be a DAG. We say that $P$ is Markovian with respect to $\mathcal{G}$ if it satisfies the following properties.
@@
These Markov properties are equivalent if the joint distribution of $V$ has a density with respect to a product measure (Lauritzen, 1996, Theorem 6.22).
@@
Different graphs may encode the same set of conditional independence relationships. For example, the graphs $A\rightarrow B\rightarrow C$ and $A\leftarrow B\rightarrow C$ both encode the assumption that $A\perp\!\!\!\perp C\mid B$. In this case, these graphs are called Markov equivalent. Graphs that are Markov equivalent cannot be distinguished based solely on conditional independencies in observational data. However, they may be distinguished in the presence of additional assumptions about the joint distribution, such as constraints on the functional form of the conditional distributions. For detailed discussions, see Example 3 in Section 4.2 and Peters et al. (2017, §4 & §7).
@@
It can be shown that the set of distributions unfaithful to a graph $\mathcal{G}$ has Lebesgue measure zero (Spirtes et al., 2000, Theorem 3.2); see also Uhler et al. (2013). This result, however, comes with an important caveat: there exist sequences of faithful distributions that converge arbitrarily close to unfaithful distributions. A more robust alternative is the concept of $\lambda$-strong faithfulness, which requires that the dependence strength (e.g., partial correlation) between d-connected variables exceeds some threshold $\lambda>0$. Unlike exact faithfulness, the set of distributions that are not $\lambda$-strong faithful may occupy a substantial volume in the space of distributions (Uhler et al., 2013).
@@
Strictly speaking, the phrase “sample uniformly” is not well defined without first specifying a parameterization, a point related to Borel’s paradox. Conditional distributions on measure zero sets can depend on the chosen parameterization (Billingsley, 1995, p. 441, Problem 33.1); see also Wang (2022). In particular, subsets such as $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ are lower dimensional manifolds within the probability simplex, so there is no canonical uniform measure on them. For $\mathcal{S}_{1}$ and the associated DAG $B\rightarrow D\leftarrow A$, we make this precise by drawing each conditional probability table independently from a $\text{Dirichlet}(1,\ldots,1)$ prior and combining them using the DAG factorization, under which the set of unfaithful distributions has Lebesgue measure zero. For $\mathcal{S}_{2}$, we describe in Appendix A1 a sampling scheme under which a distribution satisfying $A\perp B$ and $A\perp D$ almost surely does not satisfy $A\perp(B,D)$.
@@
So far, we have described how DAGs encode statistical conditional independencies via the factorization of the observed data distribution. To endow a DAG with causal meaning, we must go beyond this statistical interpretation and commit to assumptions about how the system behaves under external interventions (Pearl, 2009, Definition 1.3.1); see also Peters et al. (2017, Definition 6.32).
@@
Eqn. (6) was originally introduced by Robins (1986) and termed the g-computation algorithm formula or simply the g-formula. It was subsequently independently re-discovered and named the manipulation theorem (Spirtes et al., 2000) and the truncation formula (Pearl, 2009).
When a causal DAG is postulated, it is often assumed that the set of variables $V$ is causally sufficient, in the sense that there are no unmeasured common causes $C\notin V$ that cause two or more variables in $V$ (Spirtes, 2010). This assumption may be slightly relaxed to the notion of interventional sufficiency, which is sufficient to ensure that the truncation formula (6) holds; see Peters et al. (2017, § 9.1) for a detailed discussion. When unmeasured common causes do exist between variables, it is conventional to represent them by adding bidirected edges ($\leftrightarrow$) between the affected nodes, resulting in an Acyclic Directed Mixed Graph (ADMG) (Richardson, 2003). Formally, such graphs can be obtained as latent projections of DAGs with unobserved variables (Verma and Pearl, 1991; Richardson and Spirtes, 2002).
@@
When all variables in the DAG are observed, it is sufficient to include all the parents of $A$ in the adjustment set $L$, in which case the average causal effect (ACE) is identified by eqn. (2) (Pearl, 2009, Theorem 3.2.2). More generally, even if some variables in the DAG are unobserved, the following backdoor criterion provides a graphical condition under which causal effects can still be identified from observational data.
@@
Under this criterion, the ACE is still identified by eqn. (2) (Pearl, 2009, Theorem 3.3.2).
@@
Although Pearl (2009, § 3.2.1, eqn. (3.2)) interprets the causal semantics of a DAG using the NPSEM-IE framework (see also Section 3.2), the proof of the backdoor criterion (Pearl, 2009, Theorem 3.2.2 and § 11.3.3) relies only on the g-formula (6), and therefore holds under the causal DAG model.
The backdoor criterion can be operationalized using the following simple rules (Pearl, 2009, § 11.3.1), such that after conditioning on $L$,
@@ 2. Rule 2:
The first rule of the backdoor criterion suggests that adjusting for post-treatment variables may introduce bias. To illustrate, consider the causal DAG in Figure 2(a), where the treatment $A$ has two descendants, $S$ and $Z$, in addition to the primary variables $A$ and $Y$. Adjusting for $S$ clearly violates Rule 2 above. Now consider adjusting for $Z$. On the surface, $Z$ does not lie on any path from $A$ to $Y$, and $S$ is not a collider on the path $A\rightarrow S\rightarrow Y$. However, every observed variable has an associated error term, and in panel (a) we explicitly display $\varepsilon_{S}$, the error for $S$, since it plays a key role here. Conditioning on $Z$ induces a dependence between $A$ and $\varepsilon_{S}$, which in turn opens the non-causal path $A\leftrightarrow\varepsilon_{S}\to S\to Y$. Equivalently, this corresponds to the undirected path $A-\varepsilon_{S}-S-Y$ in the induced moral graph. Another way to see this is by marginalizing out $S$ and examining the resulting causal DAG in Figure 2(b). In this graph, it becomes clear that $Z$ is a collider on the non-causal path $A\rightarrow Z\leftarrow\varepsilon_{S}\rightarrow S$, and hence adjusting for $Z$ may introduce collider bias into the causal effect estimate (Pearl, 2009, § 11.3.1).
@@
When the backdoor criterion fails to apply directly, identification may still be possible through Pearl’s do-calculus (Pearl, 1995), a set of three transformation rules that enable the conversion of interventional expressions (e.g., $P_{\text{do}(a)}(Y)$) into estimable quantities from the observational distribution. While the backdoor criterion suffices for simple DAGs, complete graphical criteria for covariate adjustment in ancestral and partially directed graphs have been developed (Perković et al., 2018). In fact, any identifiable interventional distribution under a causal DAG model that may involve unmeasured nodes can be derived through recursive application of the do-calculus rules, as formalized by the ID algorithm (Tian and Pearl, 2002; Huang and Valtorta, 2006; Shpitser and Pearl, 2006). See Tian and Shpitser (2010) for a review of these results.
It is important to note that this result pertains to identifiability under the causal DAG model alone. Given additional assumptions, identification may still be possible even when the ID algorithm fails to apply. For instance, one can show that the ACE is not identifiable under the instrumental variable model alone (Balke and Pearl, 1997). However, it can still be identified under additional effect homogeneity assumptions (Wang and Tchetgen Tchetgen, 2018, Theorem 1); see also Hartwig et al. (2023, Theorem 1) and Dong et al. (2025, Theorem 1).
@@
There are more sophisticated algorithms that also identify counterfactual quantities, such as those discussed in Remark 3 (Shpitser and Pearl, 2007, 2008). These algorithms, unlike the ID algorithm mentioned above, rely on the NPSEM-IE framework.
@@ 3 Translating Between the Three Causal Frameworks
The question of which causal framework is best suited for practical applications has been the subject of extensive debate in the causal inference literature (e.g. Pearl, 1995; Rubin, 2004; Lauritzen, 2004; Robins and Richardson, 2011; Richardson and Robins, 2023); see also Pearl (2009, p. 106). While each approach, including potential outcomes, structural causal models, and graphical models, offers unique strengths, efforts have been made to reconcile and translate among them. Much like natural languages, certain causal concepts are more naturally expressed in one framework than another, making such translations both valuable when possible but, at times, inherently limited. In this section, we review key attempts to bridge these perspectives. Detailed comparisons of these frameworks are deferred to Section 4.
@@ 3.1 The Potential Outcomes and NPSEM-IE Frameworks
Under an NPSEM-IE, each endogenous variable $V_{j}$ is determined by a deterministic function of its parents and an exogenous noise variable $\varepsilon_{j}$, with the collection $\{\varepsilon_{j}\}$ assumed to be jointly independent. This structural system implicitly defines a rich set of potential outcomes. For example, consider the NPSEM in (4). In this case, the potential outcome $Y(l,a)$ is obtained by intervening to set $L=l$, $A=a$, and evaluating the outcome function at the realized value of the noise term $\varepsilon_{Y}$ (Pearl, 2009, §3.6.3, §7.1):
@@
see also Strotz and Wold (1960). In fact, researchers have proposed expressing NPSEMs using potential outcomes notation to clearly distinguish them from standard statistical regression models (e.g. Imbens, 2014; Richardson and Robins, 2014, 2023). For example, the NPSEM in (4) can be represented as
@@
In the NPSEM defined in (3), the set of input variables $U_{j}$ for a given node $V_{j}$ often does not include all variables that precede $V_{j}$ in the causal ordering. The omission of these variables reflects an implicit exclusion restriction on the corresponding potential outcomes (Pearl, 2009, p. 101, eqn. (3.55)):
@@
Eqn. (13) is often known as the exclusion restriction in the IV literature (Angrist et al., 1996).
@@
Robins and Richardson (2011); Richardson and Robins (2013) formalize the converse direction by showing how one can construct an NPSEM-IE from a given system of potential outcomes using what they term one-step-ahead counterfactuals, i.e., the potential outcome of a variable under interventions on all of its parents (Robins, 1986). This “counterfactual interpretation of the error terms in structural equation models” was also noted in Pearl (2009, §7.4.5) and referred to as the “canonical representation of structural causal models” by Peters et al. (2017, §3.4).
@@ 3.2 NPSEM-IE and Causal DAG
The first claim in Proposition 1 follows from Pearl (2009, Theorem 1.4.1); see also Peters et al. (2017, Proposition 6.31). The second claim holds because, under the autonomy assumption, the structural equation for the intervened variable is replaced while all others remain unchanged; see also Strotz and Wold (1960, §2), Pearl (2009, p. 72 eqn. (3.10)) and Peters et al. (2017, p. 109, eqn. (6.7)).
@@
In fact, a stronger claim holds: the Single-World Intervention Graph (SWIG) model (see §3.3 below) with interventions on all variables also implies a causal DAG model (Richardson and Robins, 2023, Theorem 11). This is stronger because the NPSEM-IE is a strict submodel of the SWIG model associated with the same DAG. The converse is not necessarily true. For example, the SWIG in Figure 3 implies the so-called extended g-formula (Robins et al., 2004) that
@@
The Functional Representation Lemma (El Gamal and Kim, 2011, p. 626, Appendix B) ensures that any random variable with a known conditional distribution can be expressed as a measurable function of its parents and an independent noise variable. A concrete construction is provided in Peters et al. (2017, Appendix C.9).
@@
where each $f_{j}$ is a measurable function, each $\varepsilon_{j}\sim\text{Uniform}[0,1]$, and the noise variables $\varepsilon_{1},\dots,\varepsilon_{p}$ are jointly independent. The NPSEM (16) with independent errors is compatible with both the causal DAG $\mathcal{G}$ and the joint distribution $P$. A similar construction for the discrete case can be found in Druzdzel and Simon (1993).